At Johnson Middle school, 170 students buy 3 ring binders from school store. This number has been increasing at a rate of about 25 per year. The number of students buying 2 ring binders is 300 and has been decreasing at a rate of about 15 per year. At these rates, when will the number of 2 ring binders sold be half the number of 3 ring binders sold?
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

170+25x=2(300-15x)
170+25x=600-30x
25x+30x=600-170
55x=430
x=430/55
x=7.818 yrs. 2 ring bindes will be half as many as the 3 ring binders.
Proof:
170+25*7.818=2(300-15*7.818)
170+195.45=2(300-117.27)
365.45=2*182.73
365.45~365.46

Answer by stanbon(75887)

At Johnson Middle school, 170 students buy 3 ring binders from school store. This number has been increasing at a rate of about 25 per year.
---
#(3 ring) = 170+25x
------------------------------------
The number of students buying 2 ring binders is 300 and has been decreasing at a rate of about 15 per year.
---
#(2 ring) = 300-15x
==============================
At these rates, when will the number of 2 ring binders sold be half the number of 3 ring binders sold?
---
Equation:
Solve: 300-15x = (1/2)(170+25x)
---
600 - 30x = 170+25x
55x = 430
x = 7.82 years
--------------------------
Cheers,
Stan H.